OpenStudy (anonymous):

Exponential Equations help? Determine the exact value of x: (1/8)^(x-3) = 2 * 16^(2x+1)

OpenStudy (anonymous):

Are you using logarithms to solve?

OpenStudy (anonymous):

can it be solved without logarithms??

OpenStudy (anonymous):

I don't know how to solve w/o logs

OpenStudy (anonymous):

okay...logarithms are fine then :)

OpenStudy (anonymous):

Haha :) First, take the log of both sides

OpenStudy (anonymous):

Then apply your log rules...in other words......

OpenStudy (anonymous):

\[\log (1/8)^{x-3}= (x-3) \log (1/8)\]

OpenStudy (anonymous):

and \[\log (2 * 16^{2x +1}) = \log 2 + \log 16^{2x+1}\]

OpenStudy (anonymous):

this equals \[\log 2 + (2x+1) \log 16\]

OpenStudy (anonymous):

so now you have\[(x-3) \log (1/8)= \log 2 + (2x + 1) \log 16\]

OpenStudy (anonymous):

evaluate the logs and solve the equation for x

OpenStudy (anonymous):

There may be another way to solve, but I'm not sure

OpenStudy (anonymous):

thanks :) ..but can you explain what you mean by evaluate the logs? i've never done them before, really. :S

OpenStudy (anonymous):

ok, forget all that....i just figured out a better way :)

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

1/8 = \[2^{-3}\]

OpenStudy (anonymous):

and 16 = \[2^{4}\]

OpenStudy (anonymous):

This gives you \[2^{-3(x-3)}= 2^{1}*2^{4(2x+1)}\]

OpenStudy (anonymous):

For the exponents, use distributive property

OpenStudy (anonymous):

\[2^{-3x+9}= 2^{1}*2^{8x +1}\]

OpenStudy (anonymous):

on the right side, you can combine those by adding expontents

OpenStudy (anonymous):

\[2^{-3x+9}=2^{8x+2}\]

OpenStudy (anonymous):

Now, since the bases are equal, this means the exonents are =, so\[-3x + 9 = 8x + 2\]

OpenStudy (anonymous):

Then, solve the equation for x

OpenStudy (anonymous):

haha....that was easier than logs :)

OpenStudy (anonymous):

yay! this makes so much sense to me. thanks so much :)

OpenStudy (anonymous):

You're very welcome!